# Extracting Probabilistic Knowledge from Large Language Models for Bayesian Network Parameterization

### v1

@@ 5 Experiments
5.1 Dataset of Eighty Bayesian Networks
Our experiments use bnRep (Leonelli, 2025), a publicly available collection of BNs spanning diverse domains such as medicine and engineering. These BNs vary in size, ranging from 5 to 50 nodes. Each BN in bnRep has an associated publication detailing its construction, data sources, and domain-specific context. Most of these papers were published between 2020 and 2024, making them well-suited for LLMs incorporating recent knowledge. Each BN is accompanied by ground-truth parameters, which enable us to evaluate our predictions. We preprocess bnRep to select only discrete BNs and eliminate BNs with missing entries in CPTs, yielding a final dataset of 80 networks. We extract node definitions and states from each BN’s repository paper. Additional details, including preprocessing steps, are provided in Appendix A.
5.2 Metrics, and Baselines
As of LLMs, we utilize GPT-4o and its mini variant (OpenAI et al., 2024) in most of our experiments, along with Claude Sonnet 3.5 (Anthropic, 2024) and Gemini-Pro 1.5 (Team et al., 2024). We evaluate the LLM’s estimated BN parameters using Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) compared to ground-truth parameters in bnRep. Specifically, we report the BN KL divergence, defined as the KL divergence computed over the BN variables’ joint distribution, evaluating the resulting BN’s overall quality. For details on the LLMs and their hyperparameters, see Appendix B, and for the explanation of KL divergence and BN KL divergence calculation, refer to Appendix C.
We evaluate our methods against multiple baselines. All outputs from these baselines are normalized as necessary to ensure valid probability distributions. These baselines are: (1) Random number generator; (2) Uniform baseline generating equal probabilities for each row of the CPT, providing basic, uninformed estimations; (2) LLM (Random) baseline involving intentionally incorrect queries, where the original variable names are randomly replaced. This is done to assess whether LLMs utilize the content of the provided questions to generate their answers. For instance, we query the LLM "What is the probability that construction time is true given that lung cancer is true?" instead of the correct question regarding smoking; (3) LLM (No Context) baseline in which queries are presented without contextual explanations, exploring scenarios where the node meanings and number of states cannot be directly inferred from the in-context information; (4) LLM (Token Probability) baseline, which directly uses the LLM’s internal probabilities assigned to tokens representing node states (e.g., probability of generating the token "True"), rather than explicitly generated numerical probabilities extracted from the model’s textual responses; (5) LLM Full Dist baseline wherein the LLM is prompted for the complete distribution of the node corresponding to a specific combination of parent states (a CPT row), and the response is received as a tuple of numbers. The order of the states is given in the prompt; and (6) MLE-# is a statistical baseline obtained by maximum likelihood estimation using data sampled from the ground-truth BN where the BN is sampled exactly # times using forward sampling. For results obtained using other sampling methods, refer to the Appendix D.
5.3 Can LLMs Estimate Probabilities Using their Internal Knowledge?
@@
To evaluate how EPK compares to other baseline models, we analyze the distributions of BN KL divergence across all 80 BNs, as depicted in Figure 8. As expected, the parameter estimation of statistical models (MLE-#) progressively improves as the size of the sample data increases. Models trained on minimal data demonstrate notably poor performance, similar to random guessing. Interestingly, the worst performing models, performing worse than random, are the “Token Probability” models. This aligns with previous research, which found that raw token probabilities from LLMs alone are insufficient for effective uncertainty/probability estimation (Xiong et al., 2024) and require additional processing steps like fine-tuning (Tao et al., 2024).
@@
The next weakest results are observed among the random generators, including GPT-4o (Random) and GPT-4o-mini (Random), next to the baselines that do not observe the correct context for nodes and their states, GPT-4o (No Context) and GPT-4o-mini (No Context). We observe that LLM (Random) models slightly surpass the outputs of the Random number generator baseline. However, these improvements only reflect the non-uniformity of random number generation by LLMs, influenced by factors such as text-generation sampling methods and model architecture choices (Hopkins et al., 2023). Of all the baseline models, the uniform predictor performs best. This result aligns with information theory, which suggests that, in the absence of knowledge, a uniform distribution naturally provides the lowest KL divergence based on uncertainty (Cover and Thomas, 2006).
@@
Lastly, both the LLM (Full Dist) baseline and EPK outperform the uniform baseline. However, while the LLM (Full Dist) baseline yields informed estimates, it consistently falls short of EPK with a higher median KL divergence and a greater standard deviation across the BNs. The MLE-# model requires at least 30 samples to match the median KL divergence of EPK. Even then, its standard deviation remains higher, indicating that EPK is more robust. Overall, these results demonstrate the capability of LLMs to provide meaningful probability distributions based on their internal knowledge.
5.4 Can LLMs’ probability estimations serve as Expert Priors?
Figure 3: Distribution of BN KL divergence across 80 BNs using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the linear pooling method. The data sample size is shown in parentheses. The MLE-# models only use data samples, and the EPK models only use LLMs for probability estimations.
@@ 5.4 Can LLMs’ probability estimations serve as Expert Priors?
We evaluate the effectiveness of EDP, combining LLM-derived distributions with empirical data, using two methods: priors as pseudocounts and linear pooling. Figure  3 displays the distributions of BN KL divergences obtained by combining GPT-4o and Uniform priors with various sample sizes of data with the linear pooling method. We chose GPT-4o because it has the best performance among the evaluated LLMs. At the same time, the Uniform model serves as the best-performing baseline in our experiments and is typically used in the absence of prior information. The method using priors as pseudocounts shows the same trend, and its results are presented in Appendix D to save space.
@@
Regardless of the combination method, GPT-4o predictions consistently outperform the Uniform model baseline, proving its use as a better prior. The advantage of incorporating GPT-4o priors is most notable at smaller sample sizes, i.e., 3 to 10 samples. The combination of even minimal data in EDP significantly outperforms EPK and the MLE model with 30 data samples. Additionally, EDP improves the median KL divergence even when abundant data, 1000 or more samples, is used. It also effectively reduces the standard deviation of the predictions, enhancing model robustness. This improvement in median KL divergence at large sample sizes occurs because nodes with unlikely parent combinations rarely receive data. This is supported by the absence of KL divergence improvement at high sample sizes with alternative methods that separately sample data for each CPT row (See Appendix D). However, these alternative, unrealistic methods still yield decreased standard deviation and improved robustness. These findings confirm that incorporating LLM predictions as expert-driven priors is beneficial and enhances performance and robustness.
@@ 6 Discussion
6.1 Quality of Individual Distributions Versus the Entire Bayesian Network
@@
In our experiments, we employed the BN KL Divergence metric to evaluate the quality of the predicted BNs. However, this metric assesses the entire BN, meaning that a few poorly predicted nodes might disproportionately influence the evaluation. To address this limitation, we also analyzed the CPT KL Divergence, which computes the average KL divergence across all individual CPT rows within each BN. This alternative measure evaluates the quality of individual distributions rather than the BN as a whole. Using CPT KL Divergence, we observed that the overall trends of our results remained consistent. Additional diagrams illustrating these findings are provided in Appendix D.
6.2 Performance Variations Among LLMs Across Different Bayesian Networks
@@
Among the evaluated LLMs, GPT-4o consistently exceeds the performance of other LLMs, though specific models perform better within specialized domains. For instance, Claude 3.5, Gemini-pro, and GPT-4o achieve the best results on BNs related to engineering, business, and medical domains, respectively. Furthermore, there are inherent differences in prediction behavior among these LLMs, likely attributed to their respective training methodologies. Specifically, Claude 3.5 performed best on BNs with low entropy probabilities, but showed the poorest performance on BNs with high entropy probabilities among the LLMs, indicating an overly confident prediction behavior. In contrast, Gemini-pro showed the opposite trend, whereas GPT-4o had a more balanced prediction profile.
6.3 Prompt Styles and Chain-of-Thought
In our experiments, consistent with previous research (Xiong et al., 2024), we observed notable volatility in token probabilities, with significant variations depending on preceding tokens. However, their behavior remained robust when prompting the LLMs to verbally generate probabilities as textual output. Experimental results remained consistent whether we instructed the model to explain its reasoning before providing probabilities or to output probabilities directly. These findings align with prior research on zero-shot regression (Nafar et al., 2025), indicating that chain-of-thought reasoning did not affect results. Furthermore, other prompting styles or additional contextual information taken from relevant Bayesian Network publications also did not influence the results. Our experiments suggest that although LLMs clearly possess and can articulate probabilistic knowledge, this capability remains consistent irrespective of explicit explanatory reasoning.
6.4 Handling Larger Parent Sets and States
@@
Intuitively, it makes sense that LLMs may struggle to provide informed predictions for more complex queries involving nodes with many parent nodes or states. We use the CPT KL Divergence metric to assess predictions among these nodes, which averages the KL divergence across all individual CPT rows rather than evaluating the entire BN. In our experiments with realistic BNs, LLM performance in these scenarios still surpassed our baseline models. LLMs consistently outperformed baselines in queries involving up to 5 parent nodes. Additionally, the LLMs performed better than baselines for nodes with 2 or 3 states. Nodes exceeding 3 states are rare in realistic BNs. Only 4 BNs had nodes restricted to 4 states, whereas 11 featured nodes with 5 states. A notable exception was the ’DustExplosion’ BN, which contained nodes with both 4 and 5 states. This network, designed to predict explosion probabilities in industrial environments, proved challenging for the LLMs. Apart from this isolated case, the LLMs consistently outperformed baseline methods. For detailed results demonstrating the performance with varying numbers of parent nodes and states, refer to Appendix E.
6.5 Trade-off Between EPK and Full Distribution Queries
@@
Our experiments show that EPK consistently outperforms directly requesting the complete distribution (Full Dist), achieving lower median KL divergence and improved stability. Although the Full Dist method is more straightforward and requires fewer queries, it leads to less precise and more variable estimates. Additionally, our proposed approach remains cost-effective by avoiding extensive reasoning or explanations, as most LLM-related costs derive from text generation. Thus, we recommend EPK for applications prioritizing accuracy and consistency, while Full Dist might be suitable in scenarios where cost saving is paramount.
6.6 Implications of the EDP method for Automated BN Construction
@@
The demonstrated capability of LLMs, particularly GPT-4o, in the EDP method has significant implications for automating BN construction. Traditionally, parameterization of Bayesian networks relies heavily on expert input, making the process labor-intensive and costly. However, cost is not the only limitation associated with expert-driven approaches. In some cases, experts may not even be available, and when they are, their expertise must be evaluated (Hald et al., 2016) or aggregated with the opinion of other experts (McAndrew et al., 2021) before their views can be utilized. Also, expert opinions are inherently susceptible to errors such as overconfidence and bias (Soll et al., 2024). Utilizing an LLM proficient across diverse domains holds enormous potential for advancing automation in BN parameterization.
6.7 Combining LLMs’ Priors with Small Data for Extraction of Probabilistic Knowledge
@@
In Section 5.4, we showed the potential of using LLM predictions in place of expert-driven priors for constructing BNs. Incorporating even a small amount of data, for example, just 3 samples, together with LLM predictions, yields a lower BN KL divergence than models trained purely on 30 data samples or those relying solely on LLM predictions. These results highlight a promising application wherein LLMs use minimal external data points to rapidly refine their predictions for probabilistic queries. Such a small amount of data could be supplied to the LLM in various ways, such as being obtained online by querying information from publicly available sources (e.g., occurrences of lung cancer among smokers). Combining LLMs’ priors with data in the EDP approach not only improves data efficiency and outperforms uniform baseline methods but also provides an exciting possibility for real-time improvement of LLM-generated probabilistic queries with minimal data input.

### v2

@@ 5 Experiments
Dataset of Eighty Bayesian Networks
Our experiments use bnRep (Leonelli 2025), a publicly available collection of BNs spanning diverse domains such as medicine and engineering. These BNs vary in size, ranging from 5 to 50 nodes. Each BN in bnRep has an associated publication detailing its construction, data sources, and domain-specific context. Most of these papers were published recently in the years between 2020 and 2024. Each BN is accompanied by ground-truth parameters, which enable us to evaluate our predictions. We preprocess bnRep to select only discrete BNs and eliminate BNs with missing entries in CPTs, yielding a final dataset of 80 networks. We extract node definitions and states from each BN’s repository paper. Additional details, including preprocessing steps, are provided in Appendix A.
Metrics and Baselines
For LLMs, we use GPT-4o and its mini variant (OpenAI et al. 2024), along with Claude 3.5 Sonnet (Anthropic 2024), Gemini-Pro 1.5 (Gemini Team et al. 2024), and DeepSeek-V3 (DeepSeek-AI et al. 2025b). We include reasoning models, o3 and its mini variant (OpenAI 2025), and DeepSeek-R1 (DeepSeek-AI et al. 2025a). We evaluate the LLM’s estimated BN parameters using Kullback–Leibler (KL) divergence (Kullback and Leibler 1951) compared to ground-truth parameters in bnRep. Specifically, we report the BN KL divergence, defined as the KL divergence computed over the BN variables’ joint distribution, evaluating the resulting BN’s overall quality. Refer to Appendix B for an overview of KL divergence, and BN KL divergence, and refer to Appendix C for details on the LLMs and hyperparameters.
We evaluate our methods against multiple baselines. All outputs from these baselines are normalized as necessary to ensure valid probability distributions. These baselines are: (1) Random number generator; (2) Uniform baseline generating equal probabilities for each row of the CPT, providing basic, uninformed estimations; (3) LLM (Random) baseline involving intentionally incorrect queries, where the original variable names are randomly replaced. This is done to assess whether LLMs utilize the content of the provided questions to generate their answers. For instance, we query the LLM ”What is the probability that construction time is true given that lung cancer is true?” instead of the correct question regarding smoking; (4) LLM (No Context) baseline in which queries are presented without contextual explanations, exploring scenarios where the node meanings and number of states cannot be directly inferred from the in-context information; (5) LLM (Token Probability) baseline, which directly uses the LLM’s internal probabilities assigned to tokens representing node states (e.g., probability of generating the token “True”), rather than explicitly generated numerical probabilities extracted from the model’s textual responses; (6) MLE-# is a statistical baseline obtained by maximum likelihood estimation using # data samples; (7) Uniform-# applies the same pseudocount updating method as EDP with # data samples, but uses a uniform prior instead of LLM-derived probabilities. Data samples are obtained from the ground-truth BN where the BN is sampled # times using forward sampling. For results obtained using other sampling methods, refer to Appendix D.
Can LLMs Estimate Probabilities Using Their Internal Knowledge?
@@
To evaluate how SepState and FullDist compare to other baseline models, we analyze the distributions of BN KL divergence across all 80 BNs, as depicted in Figure 2. The worst-performing models, which perform worse than random, are the “Token Probability” models, shown with yellow boxes. This aligns with previous research, which found that raw token probabilities from LLMs alone are insufficient for effective uncertainty/probability estimation (Xiong et al. 2024) and require additional processing steps like fine-tuning (Tao et al. 2024). The next weakest results are observed among the random generators, alongside the baselines that do not receive the context of nodes and their states. In our experiments, LLM (Random) models slightly surpass the outputs of the Random number generator baseline. However, these improvements only reflect the non-uniformity of random number generation by LLMs, influenced by factors such as text-generation sampling methods and model architecture choices (Hopkins, Renda, and Carbin 2023). Of all the baseline models, the uniform predictor performs best. This result aligns with information theory, which suggests that, in the absence of knowledge, a uniform distribution naturally provides the lowest KL divergence based on uncertainty (Cover and Thomas 2006).
@@
Both FullDist and SepState outperform the uniform baseline in all non-reasoning models. o3 and DeepSeek-R1 trail their non‑reasoning counterparts, hinting that high levels of reasoning do not translate to better probabilistic estimation. While the FullDist yields informed estimates, it consistently falls short of SepState with a higher median KL divergence and a greater standard deviation, except in the DeepSeek model. These results confirm the shortcomings of the FullDist scheme and establish SepState as a superior method to extract a full probability distribution from LLMs. Overall, these results demonstrate the capability of LLMs to provide meaningful probability distributions, laying the groundwork for treating them as informative priors in EDP.
Can LLMs’ Probability Distribution Estimates Serve as Expert Priors?
@@
In this section, we evaluate the effectiveness of EDP, which combines the LLM-derived distributions with empirical data, using priors as pseudocounts. Figure 3 displays the distributions of BN KL divergences obtained by combining various sample sizes of data with GPT-4o priors (EDP-#) and Uniform priors (Uniform-#). The Uniform prior is the conventional choice in the absence of prior information and serves as a baseline prior in our experiments.
@@
EDP predictions consistently outperform the Uniform-# baseline, proving its use as a better prior. The advantage of EDP is most notable at smaller sample sizes, i.e., 3 to 30 samples. The combination of even minimal data in EDP significantly outperforms SepState and the MLE model with 30 data samples. Additionally, EDP still improves the median KL divergence when more data are available, e.g., 1,000 samples. It also effectively reduces the standard deviation of the predictions, enhancing model robustness. This improvement in median KL divergence at large sample sizes occurs because nodes with unlikely parent combinations rarely receive data. When data are sufficiently large, such as 10k samples in our setting, MLE achieves the best KL divergence, as expected. However, this outcome relies on our forward sampling procedure, which ensures every node of the BN is sampled. In the real world, data are often sparse or biased. In this case, even with 10k samples, EDP can still provide benefits, as we will show in the next section.
For brevity, the main text reports EDP results only for GPT‑4o. The corresponding EDP plots for every other LLM priors that outperforms the Uniform baseline in Figure 2 display the exact same pattern as EDP with GPT-4o and are included in Appendix D. These findings confirm that incorporating LLM predictions as expert-driven priors is beneficial and enhances the performance and robustness of parameter estimation in Bayesian Networks.
Figure 3: Boxplot of distribution of BN KL divergence over 80 networks, contrasting models with GPT‑4o priors (EDP‑#), uniform priors (Uniform‑#), and data‑only estimates (MLE‑#). The numeral after the “‑” denotes the sample size.
What Is EDP’s Impact on Downstream Tasks?
We have shown that EDP lowers BN KL divergence, but this alone does not guarantee better performance on downstream tasks. To test whether improved probability estimates translate into downstream gains, we focus on classification, one of the most common BN applications. Following the experimental setting of (Carli et al. 2022), we select nine real‐world datasets that are routinely used for BN‐based classification studies, each with one target random variable. Every dataset is stratified into 80% training and 20% test splits. For each dataset, two BN structures are used: (i) a structure discovered via Hill Climbing, learned based on the full training data, and (ii) a simpler Naïve Bayes structure. Besides the full‐data regime, we simulate low‐resource scenarios by using the same network structure but restricting the training set to 20 and 10 samples for parameter estimation, and averaging the results over 5 random runs. We evaluate two LLMs, GPT‑4o and DeepSeek‑V3, using Chain-of-Thought (COT) (Wei et al. 2022), SepState, MLE, and EDP.
Table 1 reports macro F1 scores of these methods over 9 datasets and compares them for each network structure. EDP improves over pure data fitting (MLE), especially when data are scarce. The advantage of EDP is most evident in the derived structure from Hill Climbing because learning parameters for its complex structure is more difficult with scarce data. With full training sets, EDP still edges out or ties MLE in 12 out of 18 cases, which confirms the language models can help reduce the bias caused by sampled data. The only notable dip appears in HouseVotes84 and puffin, two datasets whose full-data MLE classifiers already touch the optimal ceiling (0.94–1.00). Even then, EDP does not significantly harm performance.
The trend holds when switching from a Hill Climbing structure to Naïve Bayes, and EDP surpasses MLE at every data size. This trend also remains consistent when we test DeepSeek-V3 (tables are moved to Appendix D for space). The only difference is that we get better classification results with EDP using the FullDist scheme for DeepSeek-V3, which achieved a better parameter estimation based on the obtained KL divergence as shown in Figure 2. Taken together, these findings corroborate the central claim that LLM-derived probability estimates are not only closer to ground truth compared to other alternative baselines but are also effective in downstream tasks.
Baseline Hill Climbing Naïve Bayes
Full Data 20 Samples 10 Samples Full Data 20 Samples 10 Samples
Dataset Size Random COT SepState MLE EDP MLE EDP MLE EDP SepState MLE EDP MLE EDP MLE EDP
HouseVotes84 435 0.52 0.50 0.24 0.94 0.91 0.87 0.80 0.78 0.85 0.40 0.94 0.91 0.92 0.90 0.91 0.87
PhDArticles 915 0.38 0.39 0.41 0.38 0.41 0.35 0.41 0.34 0.41 0.45 0.44 0.44 0.36 0.41 0.34 0.41
Pokemon 999 0.50 0.43 0.23 0.62 0.62 0.62 0.62 0.54 0.54 0.48 0.62 0.60 0.59 0.60 0.53 0.54
Titanic 32 0.51 0.71 0.42 0.42 0.42 0.42 0.42 0.42 0.42 0.57 0.11 0.55 0.22 0.57 0.42 0.56
CAD1 236 0.54 0.55 0.87 0.83 0.83 0.74 0.85 0.63 0.84 0.77 0.81 0.85 0.83 0.86 0.80 0.84
CAD2 67 0.44 0.44 0.76 0.65 0.76 0.76 0.76 0.60 0.76 0.50 0.86 0.79 0.73 0.78 0.77 0.78
Covid 10k 0.54 0.64 0.74 0.71 0.73 0.70 0.72 0.72 0.72 0.72 0.71 0.72 0.71 0.72 0.69 0.72
Puffin 69 0.50 0.43 0.63 1.00 0.93 0.99 0.91 0.97 0.91 0.78 0.93 0.85 0.90 0.88 0.87 0.88
Trajectories 10k 0.59 0.78 0.87 0.87 0.87 0.75 0.86 0.68 0.86 0.80 0.87 0.87 0.86 0.86 0.85 0.86
Average _ 0.50 0.54 0.57 0.71 0.72 0.69 0.71 0.63 0.70 0.61 0.70 0.73 0.68 0.73 0.69 0.72
Table 1: Macro‑F1 classification scores on nine datasets for BN classifiers. Columns include three methods, SepState using GPT-4o, MLE, and EDP, under two graph structures, HillClimbing and Naïve Bayes. Full Data uses the entire training split, whereas 20 and 10 Samples simulate low‑resource regimes. Within each data regime, the higher of MLE vs. EDP is bolded.
@@ 6 Discussion
Quality of Individual Distributions Versus the Entire BN. In our experiments, we used the BN KL divergence metric to evaluate the quality of the predicted BNs parameters. However, this metric assesses the entire BN, meaning that a few poorly predicted nodes might disproportionately influence the evaluation. To address this limitation, we also analyzed the CPT KL Divergence, which computes the average KL divergence across all individual CPT rows within each BN. This alternative measure evaluates the quality of individual distributions rather than the BN as a whole. Using CPT KL Divergence, we observed that the overall trends of our results remained consistent. Additional diagrams illustrating these findings are provided in Appendix D.
@@
Performance Variations Among LLMs Across Different BN Domains. Among the evaluated LLMs, GPT-4o consistently exceeds the performance of other LLMs, though specific models perform better within specialized domains. For instance, Claude 3.5, Gemini-pro, and GPT-4o achieve the best results on BNs related to engineering, business, and medical domains, respectively. Furthermore, there are inherent differences in prediction behavior among these LLMs, likely attributed to their respective training methodologies. Specifically, Claude 3.5 performed best on BNs with low entropy probabilities, but showed the poorest performance on BNs with high entropy probabilities among the LLMs, indicating an overly confident prediction behavior. In contrast, Gemini-pro showed the opposite trend, whereas GPT-4o had a more balanced prediction profile.
@@
Handling Larger Parent Sets and States. Intuitively, it is expected that LLMs may struggle to provide informed predictions for more complex queries involving nodes with many parent nodes or states. We use the CPT KL Divergence metric to assess predictions among these nodes, which averages the KL divergence across all individual CPT rows rather than evaluating the entire BN. In our experiments with realistic BNs, LLM performance in these scenarios still surpassed our baseline models. LLMs consistently outperformed baselines in queries involving up to 5 parent nodes. Additionally, the LLMs performed better than baselines for nodes with 2 or 3 states. Nodes exceeding 3 states are rare in realistic BNs. Only 4 BNs had nodes restricted to 4 states, whereas 11 featured nodes with 5 states. Within these BNs, except for the ’DustExplosion’ BN, the LLMs consistently outperformed baseline methods. ’DustExplosion’ contained nodes with both 4 and 5 states and is designed to predict explosion probabilities in industrial environments, which proved challenging for the LLMs. For detailed results demonstrating the performance with varying numbers of parent nodes and states, refer to Appendix E.
@@
Trade-off Between SepState and FullDist Schemes. Our experiments show that SepState consistently outperforms FullDist, in 7 out of 8 models, achieving lower median KL divergence and improved stability. Although the FullDist scheme is more straightforward and requires fewer queries, it leads to less precise and more variable estimates. Thus, we recommend SepState for applications prioritizing accuracy and consistency, while FullDist might be suitable in scenarios where cost saving is paramount.
@@
EDP’s Application for Automated BN Construction. The demonstrated capability of LLMs, particularly GPT-4o, in the EDP method has significant implications for automating BN construction. Traditionally, parameterization of BNs relies heavily on expert input, making the process labor-intensive, costly, and dependent on the availability of reliable experts. Utilizing an LLM proficient across diverse domains removes these barriers and holds enormous potential for advancing automation in BN parameterization.
@@
EDP with Small Data for Extraction of Probabilistic Knowledge. We showed the potential of using LLM predictions in place of expert-driven priors for constructing BNs. As shown in Figure 3, EDP with just a few data samples, such as 3, yields a lower BN KL divergence than both MLE trained on more data, like MLE-30, or the LLM-only (SepState). These results highlight a promising application wherein LLMs use minimal external data points to rapidly refine their predictions for probabilistic queries. Such a small amount of data could be supplied to the LLM in various ways, such as being obtained online by querying information from publicly available sources (e.g., occurrences of lung cancer among smokers). EDP not only improves data efficiency and outperforms uniform baseline methods but also provides an exciting possibility for real-time improvement of LLM-generated probabilistic estimates with minimal data input.